OpenStudy (anonymous):

Solve the equation. 4 times the square root of x minus 5 minus 8 = 12 A. 20 B. 30 C. 25 D. 25/4

OpenStudy (anonymous):

weird...

OpenStudy (anonymous):

4 root x - 5 - 8 = 12 4 root x = 25 root x = 100 x = 100^2 = 10000 did i get the equation wrong?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

hmmm??? how did it became 100

OpenStudy (anonymous):

root x = 25/4 x = 100/16 = 6.25

OpenStudy (anonymous):

\[4\sqrt{x-5-8}=\]

OpenStudy (anonymous):

D haha

OpenStudy (anonymous):

ah root (x-5-8) = 3 root (x-13) = 3 x - 13 = 9 x = 21

OpenStudy (anonymous):

but there is no 21 XD

OpenStudy (anonymous):

typo?

OpenStudy (anonymous):

the answer is 22 ...

OpenStudy (anonymous):

i'd make up an answer choice F. and then ask the teacher once you go to class.

OpenStudy (anonymous):

*E.

OpenStudy (anonymous):

but... 6.25 = 25/4

OpenStudy (anonymous):

but that's with the wrong initial equation, right?

OpenStudy (anonymous):

im home schooled and this is on the test i received

OpenStudy (anonymous):

hm well 20 seems like the closest answer.. according to the information in the question, the correct answer is 21.